OpenStudy (hba):

Show that the following vectors are linearly independent.(Over C or R) (1,1,1) and (0,1,-2)

OpenStudy (hba):

@Hero

Parth (parthkohli):

I am overrated. I don't know all this stuff!

OpenStudy (hba):

No worries c:

OpenStudy (hba):

@phi

OpenStudy (hba):

I also do not know this stuff,I am just starting.

OpenStudy (hba):

I know we can use wronskian,determinant and definition. Can you explain me by definition ?

OpenStudy (hba):

@zaynahf

OpenStudy (ghazi):

well for these two vectors to be linearly DEPENDENT, vector 1 should be able to be expressed as product of a constant and the another vector which means \[V _{1}=KV _{2}\] is this helpful ? or shall i explain further :)

OpenStudy (hba):

Please explain further and in depth :)

OpenStudy (ghazi):

lets say \[V _{1}= i+ j+k \] and vector \[V _{2}=j-2k\] now we can see that \[V _{1} \neq KV _{2}\] therefore these are linearly independent vectors

OpenStudy (ghazi):

if we had two vectors like \[V _{1}= i+j+k\] and \[V _{2}= 2i +2j+ 2k\] we can easily see that \[V _{1}=k V _{2}\] therefore these two are linearly dependent vectors , wherever this condition fails , it means vectors are independent

OpenStudy (hba):

Thanks a lot for helping,I got it :)

OpenStudy (ghazi):

you're welcome but do look for three vectors too , it uses determinant technique , in the above example we can see that V1=2 V2 , so its dependent

OpenStudy (hba):

@ghazi What :o Won't that be linearly independent ?

OpenStudy (ghazi):

if \[V _{1}= K V_{2}\] then its dependent because you can see that one can be expressed in terms of other but if this condition fails then both the vectors are independent

OpenStudy (hba):

I am so confused Just answer me Linearly dependent or linearly indepdent ?

OpenStudy (ghazi):

look if \[V _{1}= K V_{2}\] now , if \[V _{1}=i+j+k\] and \[V _{2}= 2i+2j+2k\] then we can see that \[V _{1}=2 *V _{2}\] therefore these vectors are LINEARLY DEPENDENT BUT IF THIS CONDITION FAILS AND ONE VECTOR IS NOT EXPRESSED BY THE HELP OF CONSTANT TO THE ANOTHER ONE THEN ITS INDEPENDENT WHICH MEANS \[V _{1} \ne K V _{2}\] as the two vectors given in your question :D

OpenStudy (hba):

Thanks a lot dude got it :)

OpenStudy (ghazi):

you're welcome :D

OpenStudy (ghazi):

if you have three vectors like \[V _{1}=i+2j+3k\]\[V _{2}= 2i+j+3k\]\[V _{3}= i+j+k\] then you have to take determinant of these p elements that is the coefficients of i, j, k in all the three vectors  if the determinant is zero then vectors are linearly independent